# Generating Multi-Table Time Series EHR from Latent Space with Minimal Preprocessing

### v1

@@ A.2 Feature count calculation
Possible Feature $\text{Possible Feature Count}=\sum_{\epsilon\in\mathcal{E}}\Bigg{(}\sum_{i\in I_{\epsilon}}\Big{(}\prod_{c\in C_{\epsilon,\text{cat}}}|V_{i,c}|\Big{)}\times\left(|C_{\epsilon,\text{num}}|+1\right)\Bigg{)},$
@@
Actual Feature $\text{Actual Feature Count}=\sum_{\epsilon\in\mathcal{E}}\Bigg{(}\sum_{i\in I_{\epsilon}}\left|C_{\epsilon,\text{cat}}^{\text{unique}}(i)\right|\times\left(|C_{\epsilon,\text{num}}|+1\right)\Bigg{)},$
@@ C.1 Single-table evaluation
  • Item-specific Pairwise Column Correlation (I-PCC): I-PCC extends PCC by computing correlation matrices for each item-specific subset of the data (e.g., all records for a specific itemid). The $\mu_{\mathrm{abs}}$ is calculated for each item’s matrix and averaged across items. As with I-CDE, items with fewer than 0.1% of records are excluded.
@@ C.2 Time-series multi-table evaluation
  • Clinical Utility: To assess clinical utility, we evaluate the performance of synthetic data on downstream predictive tasks. We defined 11 clinical prediction tasks (e.g., predicting creatinine and hemoglobin levels), following the task definitions from genhpf . Details are provided in Table 10. To address variability in the number and types of events across patients, we employ two representation methods, GenHPF and MEDS-TAB, as detailed in E.2.1. For each task, the synthetic data is split into two segments over an observation window of length $T$ (e.g., $T=12$ hours. The first half (e.g., the first 6 hours) serves as input for the predictive model, with predictions made at the midpoint ($T/2$). For multi-class tasks, the label is the last value of the lab measurement in the second half (e.g., the subsequent 6 hours). For binary tasks, the label indicates whether a medication event occurs in the second half. Models are trained on synthetic data and tested on real data, with performance measured using the Area Under the Receiver Operating Characteristic Curve (AUROC). Higher AUROC scores indicate better utility for clinical applications.
@@
  • Next Event Prediction: Next Event Prediction evaluates the synthetic data’s ability to capture temporal sequence dynamics. An LSTM-based model (lstm, ) predicts the next event’s item or drug name, formulated as a multi-label classification task to handle concurrent events. The model is trained on synthetic data and evaluated on real test data using the F1 score, which balances precision and recall for multi-label predictions. Higher F1 scores indicate better sequence modeling.
@@ D.1 Architecture details
The event compression module transforms serialized clinical event text into compact latent representations, leveraging a Vector Quantized Variational AutoEncoder (VQ-VAE) vqvae or Residual Quantization (RQ) framework rqvae , both implemented using 1D convolutional neural networks (CNNs). The processing pipeline begins with the construction of input embeddings, followed by encoding, quantization, and decoding stages.
The input embedding, denoted $\mathbf{x}_{i}^{p}\in\mathbb{R}^{L\times F}$, integrates three components adapted from an established text-based EHR prediction framework, GenHPF, to ensure robust representation of clinical data. The textual embedding ($\mathbf{x}_{\text{text},i}^{p}$) employs framework-derived embeddings for serialized event text, such as “lab item Glucose value 95 uom mg/dL”, tokenized using a Bio+Clinical BERT tokenizer bioclinicalbert . The type embedding ($\mathbf{x}_{\text{type},i}^{p}$) assigns categorical labels to tokens, distinguishing table names, column names, and values to capture the relational structure of the data. The digit-place embedding ($\mathbf{x}_{\text{dpe},i}^{p}$) enhances numeric tokens by tokenizing digits with spaces (e.g., “123.45” as “1 2 3 . 4 5”) and assigning position-specific types for integer and decimal places. These components are combined as $\mathbf{x}_{i}^{p}=\mathbf{x}_{\text{text},i}^{p}+\mathbf{x}_{\text{type},i}^{p}+\mathbf{x}_{\text{dpe},i}^{p}$. The main text adopts the notation $\mathbf{x}_{i}^{p}=\mathbf{x}_{\text{text},i}^{p}$ for simplicity, though the full embedding incorporates type and digit-place embeddings to enhance structural and numeric fidelity.
@@ D.2 Postprocessing
This appendix elaborates on the postprocessing steps applied to convert synthetic patient trajectories into relational tables, aligning with the structural and semantic consistency of real data. The core steps of event-level verification and patient-level validation, which ensure valid event formats and temporal consistency, are implemented in Algorithm 1. Here, we detail the additional postprocessing steps that enforce column-specific constraints, as referenced in Section 3.4, to refine the synthetic tables’ statistical fidelity.
@@
Input: Text-based events $E$, timestamps $\mathcal{T}$, valid column names $\mathcal{C}$, valid table names $\mathcal{E}$, observation window $\mathcal{T}_{\text{max}}$
@@
Initialize $\mathcal{R}\leftarrow\emptyset$ (dictionary of tables)
for each event $e_{i}\in E$ with timestamp $t_{i}\in\mathcal{T}$ do
       Parse $e_{i}$ into table name $\tau$ and column-value pairs $\{(c_{i,j},v_{i,j})\}$
@@
      if $e_{i}$ does not conform to column-value pair format then
             Discard $e_{i}$
@@
      for each $(c_{i,j},v_{i,j})$ in $\{(c_{i,j},v_{i,j})\}$ do
             if $c_{i,j}\notin\mathcal{C}$ then
                   $c_{i,j}\leftarrow\arg\min_{c\in\mathcal{C}}\text{Levenshtein}(c_{i,j},c)$
@@
            if $c_{i,j}$ is numeric and $v_{i,j}$ contains non-numeric characters then
                   $v_{i,j}\leftarrow\text{RemoveNonNumeric}(v_{i,j})$
@@
      Add $(t_{i},\{(c_{i,j},v_{i,j})\})$ to $\mathcal{R}[\tau]$
@@
for each table $\mathcal{R}[\tau]$ in $\mathcal{R}$ do
       Sort events by timestamp $t_{i}$ in ascending order
      for each event $(t_{i},\{(c_{i,j},v_{i,j})\})$ in $\mathcal{R}[\tau]$ do
             if $i>1$ and $t_{i}<t_{i-1}$ or $t_{i}>\mathcal{T}_{\text{max}}$ then
                   Discard $(t_{i},\{(c_{i,j},v_{i,j})\})$ and all events $(t_{k},\{(c_{k,j},v_{k,j})\})$ for $k\geq i$
@@
Appendix E Training details and hyperparameters
E.1 Generative models
  • RealTabFormer realtabformer : Text-based autoregressive model for generating relational database. This model involves fine-tuning the Meta-LLaMA-3.1-8B model and a generation step using top-p sampling.
    • Fine-tuning: The Meta-LLaMA-3.1-8B model dubey2024llama was fine-tuned using QLoRA dettmers2024qlora (rank $r$=128, LoRA scaling factor $\alpha$=128, LoRA dropout=0.05) and Flash Attention-2 for memory optimization. Training used a maximum sequence length of 11,240 tokens, matching the input data. The Adafactor optimizer was applied with a learning rate of 2e-4, weight decay of 0.001, a batch size of 1, gradient accumulation (step size=4), and gradient checkpointing. Training ran for 2 epochs on a single RTX 3090 GPU, taking approximately 10 days for MIMIC-IV and 20 days for eICU.
  • SDV sdv : The Synthetic Data Vault (SDV) uses the Hierarchical ML Algorithm (HMA) Synthesizer to generate synthetic relational data, modeling individual tables with Gaussian Copulas while preserving parent-child relationships.The SDV framework’s default hyperparameters are applied for training.
  • RC-TGANrctgan A GAN-based model designed to generate synthetic relational databases by modeling parent-child relationships.
  • ClavaDDPMclavaddpm A diffusion-based model for multi-table data generation, Cluster Latent Variable guided Denoising Diffusion Probabilistic Models (ClavaDDPM) leverage Denoising Diffusion Probabilistic Models (DDPMs) to model complex tabular data distributions. By using clustering labels as intermediaries, ClavaDDPM captures long-range dependencies across interconnected tables, particularly through foreign key constraints.
@@ E.2.1 Clinical utility
  • GenHPF genhpf : GenHPF addresses the heterogeneity of multi-table EHR data by transforming all patient events into a single hierarchical textual sequence. This method sequentially lists events such as medications, infusions, and lab results in chronological order, requiring minimal preprocessing. GenHPF enables predictive models to capture the full complexity of patient records, making it suitable for multi-task learning without extensive domain-specific feature engineering. Its flexibility supports applications across diverse EHR systems, enhancing model generalizability for downstream predictive tasks.
@@
  • MEDS-TAB medstab : MEDS-TAB is an automated tabularization tool that converts Medical Event Data Standard (MEDS)-formatted EHR data into standardized tabular representations by aggregating events into fixed time intervals. It processes irregularly sampled time-series data, supporting various aggregation functions (e.g., sum, count, mean) and window sizes to handle diverse data types, including static codes, numerical values, and time-series events. With high scalability and efficiency—demonstrated by processing 500 patients from MIMIC-IV in 16 seconds with 1,410MB memory usage—MEDS-TAB simplifies predictive modeling by providing structured inputs compatible with machine learning algorithms like XGBoost, achieving competitive performance in clinical tasks.
@@ F.1 Column-wise fidelity metrics
L/itemid C 0.06 - 0.00 0.00 66.08 ±$\pm 24.72$ 78.87 ±$\pm 23.05$
L/value C 0.07 0.14 5.78 4.51 0.00 ±$\pm 0.00$ 0.77 ±$\pm 0.02$
L/valueuom C 0.04 0.03 13.78 12.49 0.07 ±$\pm 0.00$ 2.50 ±$\pm 0.02$
@@
L/priority C 0.03 0.04 27.34 29.48 44.16 ±$\pm 0.08$ 44.32 ±$\pm 0.19$
L/comments C 0.11 0.22 83.41 86.15 57.34 ±$\pm 0.61$ 63.52 ±$\pm 2.04$
L/valuenum N 0.01 0.06 11.13 9.46 116.64 ±$\pm 0.76$ 86.37 ±$\pm 1.99$
L/ref_range_lower N 0.02 0.01 19.04 17.15 60.86 ±$\pm 0.21$ 63.10 ±$\pm 0.20$
L/ref_range_upper N 0.01 0.01 19.04 17.15 30.45 ±$\pm 0.66$ 30.01 ±$\pm 0.36$
I/itemid C 0.09 - 0.00 0.00 28.17 ±$\pm 3.03$ 29.19 ±$\pm 0.13$
I/amountuom C 0.02 0.02 0.00 0.00 0.07 ±$\pm 0.00$ 0.13 ±$\pm 0.00$
I/rateuom C 0.05 0.05 37.49 39.54 0.02 ±$\pm 0.00$ 0.07 ±$\pm 0.00$
I/ordercategoryname C 0.03 0.03 0.00 0.00 0.16 ±$\pm 0.00$ 0.19 ±$\pm 0.00$
I/secondaryordercategoryname C 0.02 0.01 25.82 26.57 0.00 ±$\pm 0.00$ 0.00 ±$\pm 0.00$
I/ordercomponenttypedescription C 0.01 0.01 0.00 0.00 0.00 ±$\pm 0.00$ 0.01 ±$\pm 0.00$
I/ordercategorydescription C 0.02 0.03 0.00 0.00 0.00 ±$\pm 0.00$ 0.00 ±$\pm 0.00$
@@
I/isopenbag C 0.01 0.01 0.00 0.00 0.00 ±$\pm 0.00$ 0.00 ±$\pm 0.00$
I/amount N 0.03 0.08 0.00 0.05 88.26 ±$\pm 0.87$ 87.75 ±$\pm 2.87$
I/rate N 0.04 0.13 37.49 39.54 36.47 ±$\pm 1.69$ 62.34 ±$\pm 1.11$
I/patientweight N 0.09 - 0.00 0.00 21.23 ±$\pm 0.01$ 21.78 ±$\pm 0.02$
I/totalamount N 0.02 0.07 12.38 13.63 6.77 ±$\pm 0.08$ 8.81 ±$\pm 0.02$
I/originalamount N 0.03 0.12 0.17 0.10 28.00 ±$\pm 1.23$ 30.78 ±$\pm 0.34$
I/originalrate N 0.03 0.09 5.83 4.46 41.72 ±$\pm 3.88$ 48.13 ±$\pm 2.02$
M/drug C 0.08 - 0.00 0.00 31.00 ±$\pm 0.38$ 40.53 ±$\pm 1.01$
M/drug_type C 0.00 0.00 0.00 0.00 0.82 ±$\pm 0.07$ 1.41 ±$\pm 0.04$
M/prod_strength C 0.08 0.05 0.09 0.04 17.69 ±$\pm 0.63$ 37.79 ±$\pm 0.43$
M/dose_unit_rx C 0.03 0.02 0.09 0.04 5.93 ±$\pm 0.11$ 13.06 ±$\pm 0.28$
M/form_unit_disp C 0.03 0.04 0.09 0.05 8.16 ±$\pm 0.21$ 21.07 ±$\pm 0.45$
M/route C 0.04 0.04 0.09 0.04 31.99 ±$\pm 0.49$ 40.59 ±$\pm 0.30$
M/doses_per_24_hrs N 0.02 0.06 59.97 61.97 60.30 ±$\pm 0.11$ 60.36 ±$\pm 0.12$
M/dose_val_rx N 0.02 0.04 6.13 6.38 109.53 ±$\pm 0.20$ 105.76 ±$\pm 0.57$
M/form_val_disp N 0.02 0.03 5.59 5.99 129.93 ±$\pm 13.27$ 82.49 ±$\pm 4.43$
@@
L/labname C 0.04 - 0.00 0.00 52.57 ±$\pm 3.92$ 53.10 ±$\pm 12.26$
L/labmeasurenamesystem C 0.03 0.07 5.24 4.82 0.07 ±$\pm 0.02$ 0.59 ±$\pm 0.07$
L/labmeasurenameinterface C 0.04 0.14 6.37 5.59 54.97 ±$\pm 6.30$ 49.94 ±$\pm 0.49$
L/labtypeid C 0.03 0.06 0.00 0.00 0.00 ±$\pm 0.00$ 0.28 ±$\pm 0.04$
L/labresult N 0.01 0.11 0.93 0.64 14.68 ±$\pm 0.55$ 24.54 ±$\pm 0.65$
L/labresulttext N 0.01 0.12 0.75 0.54 19.72 ±$\pm 0.46$ 23.40 ±$\pm 0.35$
I/drugname C 0.16 - 0.00 0.00 57.38 ±$\pm 0.32$ 60.48 ±$\pm 0.20$
I/infusionrate N 0.03 0.43 55.54 63.14 77.70 ±$\pm 0.27$ 77.95 ±$\pm 1.08$
I/drugamount N 0.06 0.32 67.48 74.43 71.46 ±$\pm 1.57$ 69.29 ±$\pm 0.18$
I/volumeoffluid N 0.03 0.23 67.47 74.37 14.80 ±$\pm 0.47$ 18.11 ±$\pm 0.18$
I/patientweight N 0.03 - 90.91 95.91 12.49 ±$\pm 0.20$ 33.41 ±$\pm 0.44$
I/drugrate N 0.03 0.36 0.46 0.55 93.12 ±$\pm 0.66$ 92.25 ±$\pm 0.25$
M/drugname C 0.10 - 32.53 29.65 73.75 ±$\pm 0.17$ 81.10 ±$\pm 0.47$
M/drugivadmixture C 0.02 0.03 0.00 0.00 14.49 ±$\pm 0.07$ 16.36 ±$\pm 0.01$
M/dosage C 0.07 0.15 11.13 11.38 80.19 ±$\pm 0.08$ 81.74 ±$\pm 0.17$
M/routeadmin C 0.07 0.11 0.00 0.01 55.23 ±$\pm 0.15$ 57.59 ±$\pm 0.09$
M/frequency C 0.08 0.22 12.60 11.31 85.52 ±$\pm 0.19$ 88.22 ±$\pm 0.10$
M/prn C 0.00 0.06 0.01 0.08 10.40 ±$\pm 0.12$ 11.97 ±$\pm 0.10$

### v2

@@ A.2 Feature count calculation
$\text{Possible Feature Count}=\sum_{\epsilon\in\mathcal{E}}\Bigg(\sum_{i\in I_{\epsilon}}\Big(\prod_{c\in C_{\epsilon,\text{cat}}}|V_{i,c}|\Big)\times\left(|C_{\epsilon,\text{num}}|+1\right)\Bigg),$
@@
$\text{Actual Feature Count}=\sum_{\epsilon\in\mathcal{E}}\Bigg(\sum_{i\in I_{\epsilon}}\left|C_{\epsilon,\text{cat}}^{\text{unique}}(i)\right|\times\left(|C_{\epsilon,\text{num}}|+1\right)\Bigg),$
@@ C.1 Single-table evaluation
  • Item-specific Pairwise Column Correlation (I-PCC): I-PCC extends PCC by computing correlation matrices for each item-specific subset of the data (e.g., all records for a specific itemid). The $\mu_{\mathrm{abs}}$ is calculated for each item’s matrix and averaged across items. Items with fewer than 100 records are excluded.
@@ C.2 Time-series multi-table evaluation
  • Clinical Utility: To assess clinical utility, we evaluate the performance of synthetic data on downstream predictive tasks. We defined 11 clinical prediction tasks (e.g., predicting creatinine and hemoglobin levels), following the task definitions from Hur et al. (2023). Details are provided in Table 10. To address variability in the number and types of events across patients, we employ two representation methods, GenHPF and MEDS-TAB, as detailed in E.2.1. For each task, the synthetic data is split into two segments over an observation window of length $T$ (e.g., $T=12$ hours. The first half (e.g., the first 6 hours) serves as input for the predictive model, with predictions made at the midpoint ($T/2$). For multi-class tasks, the label is the last value of the lab measurement in the second half (e.g., the subsequent 6 hours). For binary tasks, the label indicates whether a medication event occurs in the second half. Models are trained on synthetic data and tested on real data, with performance measured using the Area Under the Receiver Operating Characteristic Curve (AUROC). Higher AUROC scores indicate better utility for clinical applications.
@@
  • Next Event Prediction: Next Event Prediction evaluates the synthetic data’s ability to capture temporal sequence dynamics. An LSTM-based model (Hochreiter and Schmidhuber, 1997) predicts the next event’s item or drug name, formulated as a multi-label classification task to handle concurrent events. The model is trained on synthetic data and evaluated on real test data using the F1 score, which balances precision and recall for multi-label predictions. Higher F1 scores indicate better sequence modeling.
@@ D.1 Architecture details
The event compression module transforms serialized clinical event text into compact latent representations, leveraging a Vector Quantized Variational AutoEncoder (VQ-VAE) Van Den Oord et al. (2017) or Residual Quantization (RQ) framework Lee et al. (2022), both implemented using 1D convolutional neural networks (CNNs). The processing pipeline begins with the construction of input embeddings, followed by encoding, quantization, and decoding stages.
The input embedding, denoted $\mathbf{x}_{i}^{p}\in\mathbb{R}^{L\times F}$, integrates three components adapted from an established text-based EHR prediction framework, GenHPF, to ensure robust representation of clinical data. The textual embedding ($\mathbf{x}_{\text{text},i}^{p}$) employs framework-derived embeddings for serialized event text, such as “lab item Glucose value 95 uom mg/dL”, tokenized using a Bio+Clinical BERT tokenizer Alsentzer et al. (2019). The type embedding ($\mathbf{x}_{\text{type},i}^{p}$) assigns categorical labels to tokens, distinguishing table names, column names, and values to capture the relational structure of the data. The digit-place embedding ($\mathbf{x}_{\text{dpe},i}^{p}$) enhances numeric tokens by tokenizing digits with spaces (e.g., “123.45” as “1 2 3 . 4 5”) and assigning position-specific types for integer and decimal places. These components are combined as $\mathbf{x}_{i}^{p}=\mathbf{x}_{\text{text},i}^{p}+\mathbf{x}_{\text{type},i}^{p}+\mathbf{x}_{\text{dpe},i}^{p}$. The main text adopts the notation $\mathbf{x}_{i}^{p}=\mathbf{x}_{\text{text},i}^{p}$ for simplicity, though the full embedding incorporates type and digit-place embeddings to enhance structural and numeric fidelity.
@@ D.2 Postprocessing
This appendix elaborates on the two-stage postprocessing pipeline applied to convert synthetic patient trajectories into relational tables, aligning with the structural and semantic consistency of real data. The first stage, converting text to tabular format, involves the core steps of event-level verification and patient-level validation (Algorithm 1). The second stage, enhancing data quality, applies additional column-specific constraints to refine the synthetic tables’ statistical fidelity, as referenced in Section 3.4.
Column-Specific Constraint Enforcement
@@
Modification and Rejection Rate Analysis
To quantify the impact of our validation protocols, we analyzed the modification and rejection rates across both postprocessing stages for the eICU and MIMIC-IV datasets.
During the initial stage, converting text to tabular format, $0.42\%$ of events were modified and $0.07\%$ were rejected in eICU. For MIMIC-IV, these rates were $0.73\%$ and $0.17\%$, respectively. At the patient level, however, rejection rates were notably higher, reaching $2.71\%$ for eICU and $18.23\%$ for MIMIC-IV.
In the subsequent stage, enhancing data quality, stricter column-specific constraints were enforced. This led to event-level rates of $1.39\%$ modified and $0.49\%$ rejected for eICU, and $1.07\%$ modified and $0.79\%$ rejected for MIMIC-IV. These stringent checks led to a substantial increase in patient-level rejection rates, which rose to $16.90\%$ in eICU and $54.53\%$ in MIMIC-IV.
The pronounced disparity between patient-level and event-level rejection rates is a direct consequence of our stringent data integrity policy. A single erroneous event necessitates the rejection of the entire patient trajectory. This effect is particularly notable for datasets with long sequences, (e.g., with hundreds of events), all of which must be accurate for the trajectory to be retained. We emphasize that this full post-processing is crucial for RawMed’s data integrity. We do not recommend partial post-processing, as it fundamentally compromises the quality of the synthetic data.
Appendix E Training details and hyperparameters
E.1 Generative models
  • RealTabFormer Solatorio and Dupriez (2023): Text-based autoregressive model for generating relational database. This model involves fine-tuning the Meta-LLaMA-3.1-8B model and a generation step using top-p sampling.
    • Fine-tuning: The Meta-LLaMA-3.1-8B model Dubey et al. (2024) was fine-tuned using QLoRA Dettmers et al. (2024) (rank rr=128, LoRA scaling factor $\alpha$=128, LoRA dropout=0.05) and Flash Attention-2 for memory optimization. Training used a maximum sequence length of 11,240 tokens, matching the input data. The Adafactor optimizer was applied with a learning rate of 2e-4, weight decay of 0.001, a batch size of 1, gradient accumulation (step size=4), and gradient checkpointing. Training ran for 2 epochs on a single RTX 3090 GPU, taking approximately 10 days for MIMIC-IV and 20 days for eICU.
  • SDV Patki et al. (2016): The Synthetic Data Vault (SDV) uses the Hierarchical ML Algorithm (HMA) Synthesizer to generate synthetic relational data, modeling individual tables with Gaussian Copulas while preserving parent-child relationships.The SDV framework’s default hyperparameters are applied for training.
  • RC-TGANGueye et al. (2023) A GAN-based model designed to generate synthetic relational databases by modeling parent-child relationships.
  • ClavaDDPMPang et al. (2024) A diffusion-based model for multi-table data generation, Cluster Latent Variable guided Denoising Diffusion Probabilistic Models (ClavaDDPM) leverage Denoising Diffusion Probabilistic Models (DDPMs) to model complex tabular data distributions. By using clustering labels as intermediaries, ClavaDDPM captures long-range dependencies across interconnected tables, particularly through foreign key constraints.
Input: Set of patients $\mathcal{P}$, Text-based events $E$, timestamps $\mathcal{T}$, valid column names $\mathcal{C}$, valid table names $\mathcal{E}$, observation window $\mathcal{T}_{\text{max}}$
@@ E.1 Generative models
Initialize $\mathcal{R}\leftarrow\text{dict of }\tau\to\text{list}$ (a dictionary mapping table names to lists of event records)
Initialize $\mathcal{P}_{\text{valid}}\leftarrow\emptyset$
for each patient $p\in\mathcal{P}$ do
   Initialize $\mathcal{R}^{\prime}\leftarrow\text{dict of }\tau\to\text{list}$ (temporary tables for patient $p$)
   Initialize $\text{valid\_event\_count}\leftarrow 0$
 for each event $e_{i}^{p}\in E^{p}$ with timestamp $t_{i}^{p}\in\mathcal{T}^{p}$ do
      Parse $e_{i}^{p}$ into table name $\tau$ and column-value pairs $\{(c_{i,j}^{p},v_{i,j}^{p})\}$
@@
    if $e_{i}^{p}$ does not conform to column-value pair format then
@@
    for each $(c_{i,j}^{p},v_{i,j^{p}})$ in $\{(c_{i,j}^{p},v_{i,j}^{p})\}$ do
       if $c_{i,j}^{p}\notin\mathcal{C}$ then
          $c_{i,j}^{p}\leftarrow\arg\min_{c\in\mathcal{C}}\text{Levenshtein}(c_{i,j}^{p},c)$
@@
       if $c_{i,j}^{p}$ is numeric column and $v_{i,j}^{p}$ contains non-numeric characters then
          $v_{i,j}^{p}\leftarrow\text{RemoveNonNumeric}(v_{i,j}^{p})$
          if $v_{i,j}^{p}$ is not numeric then
               Continue
            end if
@@
     Add $(t_{i}^{p},\{(c_{i,j}^{p},v_{i,j}^{p})\})$ to $\mathcal{R}^{\prime}[\tau]$
    $\text{valid\_event\_count}\leftarrow\text{valid\_event\_count}+1$
@@
 if valid_event_count == len($E^{p}$) then
      Add $p$ to $\mathcal{P}_{\text{valid}}$
     Add $\mathcal{R}^{\prime}[\tau]$ to $\mathcal{R}[\tau]$
   end if
end for
Sort $\mathcal{R}$ by timestamp in ascending order
for each patient $p$ in $\mathcal{P}_{\text{valid}}$ do
 for each event $(t_{i}^{p},\{(c_{i,j}^{p},v_{i,j}^{p})\})$ in $\mathcal{R}$ do
    if $i>1$ and $t_{i}^{p}<t_{i-1}^{p}$ or $t_{i}^{p}>\mathcal{T}_{\text{max}}$ then
         Discard $(t_{i}^{p},\{(c_{i,j}^{p},v_{i,j}^{p})\})$ and all events $(t_{k}^{p},\{(c_{k,j}^{p},v_{k,j}^{p})\})$ for $k\geq i$ from $\mathcal{R}$
@@ E.2.1 Clinical utility
  • GenHPF Hur et al. (2023): GenHPF addresses the heterogeneity of multi-table EHR data by transforming all patient events into a single hierarchical textual sequence. This method sequentially lists events such as medications, infusions, and lab results in chronological order, requiring minimal preprocessing. GenHPF enables predictive models to capture the full complexity of patient records, making it suitable for multi-task learning without extensive domain-specific feature engineering. Its flexibility supports applications across diverse EHR systems, enhancing model generalizability for downstream predictive tasks.
@@
  • MEDS-TAB Oufattole et al. (2024): MEDS-TAB is an automated tabularization tool that converts Medical Event Data Standard (MEDS)-formatted EHR data into standardized tabular representations by aggregating events into fixed time intervals. It processes irregularly sampled time-series data, supporting various aggregation functions (e.g., sum, count, mean) and window sizes to handle diverse data types, including static codes, numerical values, and time-series events. With high scalability and efficiency—demonstrated by processing 500 patients from MIMIC-IV in 16 seconds with 1,410MB memory usage—MEDS-TAB simplifies predictive modeling by providing structured inputs compatible with machine learning algorithms like XGBoost, achieving competitive performance in clinical tasks.
@@ F.1 Column-wise fidelity metrics
L/itemid C 0.06 - 0.00 0.00 66.08 $\pm 24.72$ 78.87 $\pm 23.05$
L/value C 0.07 0.14 5.78 4.51 0.00 $\pm 0.00$ 0.77 $\pm 0.02$
L/valueuom C 0.04 0.03 13.78 12.49 0.07 $\pm 0.00$ 2.50 $\pm 0.02$
@@
L/priority C 0.03 0.04 27.34 29.48 44.16 $\pm 0.08$ 44.32 $\pm 0.19$
L/comments C 0.11 0.22 83.41 86.15 57.34 $\pm 0.61$ 63.52 $\pm 2.04$
L/valuenum N 0.01 0.06 11.13 9.46 116.64 $\pm 0.76$ 86.37 $\pm 1.99$
L/ref_range_lower N 0.02 0.01 19.04 17.15 60.86 $\pm 0.21$ 63.10 $\pm 0.20$
L/ref_range_upper N 0.01 0.01 19.04 17.15 30.45 $\pm 0.66$ 30.01 $\pm 0.36$
I/itemid C 0.09 - 0.00 0.00 28.17 $\pm 3.03$ 29.19 $\pm 0.13$
I/amountuom C 0.02 0.02 0.00 0.00 0.07 $\pm 0.00$ 0.13 $\pm 0.00$
I/rateuom C 0.05 0.05 37.49 39.54 0.02 $\pm 0.00$ 0.07 $\pm 0.00$
I/ordercategoryname C 0.03 0.03 0.00 0.00 0.16 $\pm 0.00$ 0.19 $\pm 0.00$
I/secondaryordercategoryname C 0.02 0.01 25.82 26.57 0.00 $\pm 0.00$ 0.00 $\pm 0.00$
I/ordercomponenttypedescription C 0.01 0.01 0.00 0.00 0.00 $\pm 0.00$ 0.01 $\pm 0.00$
I/ordercategorydescription C 0.02 0.03 0.00 0.00 0.00 $\pm 0.00$ 0.00 $\pm 0.00$
@@
I/isopenbag C 0.01 0.01 0.00 0.00 0.00 $\pm 0.00$ 0.00 $\pm 0.00$
I/amount N 0.03 0.08 0.00 0.05 88.26 $\pm 0.87$ 87.75 $\pm 2.87$
I/rate N 0.04 0.13 37.49 39.54 36.47 $\pm 1.69$ 62.34 $\pm 1.11$
I/patientweight N 0.09 - 0.00 0.00 21.23 $\pm 0.01$ 21.78 $\pm 0.02$
I/totalamount N 0.02 0.07 12.38 13.63 6.77 $\pm 0.08$ 8.81 $\pm 0.02$
I/originalamount N 0.03 0.12 0.17 0.10 28.00 $\pm 1.23$ 30.78 $\pm 0.34$
I/originalrate N 0.03 0.09 5.83 4.46 41.72 $\pm 3.88$ 48.13 $\pm 2.02$
M/drug C 0.08 - 0.00 0.00 31.00 $\pm 0.38$ 40.53 $\pm 1.01$
M/drug_type C 0.00 0.00 0.00 0.00 0.82 $\pm 0.07$ 1.41 $\pm 0.04$
M/prod_strength C 0.08 0.05 0.09 0.04 17.69 $\pm 0.63$ 37.79 $\pm 0.43$
M/dose_unit_rx C 0.03 0.02 0.09 0.04 5.93 $\pm 0.11$ 13.06 $\pm 0.28$
M/form_unit_disp C 0.03 0.04 0.09 0.05 8.16 $\pm 0.21$ 21.07 $\pm 0.45$
M/route C 0.04 0.04 0.09 0.04 31.99 $\pm 0.49$ 40.59 $\pm 0.30$
M/doses_per_24_hrs N 0.02 0.06 59.97 61.97 60.30 $\pm 0.11$ 60.36 $\pm 0.12$
M/dose_val_rx N 0.02 0.04 6.13 6.38 109.53 $\pm 0.20$ 105.76 $\pm 0.57$
M/form_val_disp N 0.02 0.03 5.59 5.99 129.93 $\pm 13.27$ 82.49 $\pm 4.43$
@@
L/labname C 0.04 - 0.00 0.00 52.57 $\pm 3.92$ 53.10 $\pm 12.26$
L/labmeasurenamesystem C 0.03 0.07 5.24 4.82 0.07 $\pm 0.02$ 0.59 $\pm 0.07$
L/labmeasurenameinterface C 0.04 0.14 6.37 5.59 54.97 $\pm 6.30$ 49.94 $\pm 0.49$
L/labtypeid C 0.03 0.06 0.00 0.00 0.00 $\pm 0.00$ 0.28 $\pm 0.04$
L/labresult N 0.01 0.11 0.93 0.64 14.68 $\pm 0.55$ 24.54 $\pm 0.65$
L/labresulttext N 0.01 0.12 0.75 0.54 19.72 $\pm 0.46$ 23.40 $\pm 0.35$
I/drugname C 0.16 - 0.00 0.00 57.38 $\pm 0.32$ 60.48 $\pm 0.20$
I/infusionrate N 0.03 0.43 55.54 63.14 77.70 $\pm 0.27$ 77.95 $\pm 1.08$
I/drugamount N 0.06 0.32 67.48 74.43 71.46 $\pm 1.57$ 69.29 $\pm 0.18$
I/volumeoffluid N 0.03 0.23 67.47 74.37 14.80 $\pm 0.47$ 18.11 $\pm 0.18$
I/patientweight N 0.03 - 90.91 95.91 12.49 $\pm 0.20$ 33.41 $\pm 0.44$
I/drugrate N 0.03 0.36 0.46 0.55 93.12 $\pm 0.66$ 92.25 $\pm 0.25$
M/drugname C 0.10 - 32.53 29.65 73.75 $\pm 0.17$ 81.10 $\pm 0.47$
M/drugivadmixture C 0.02 0.03 0.00 0.00 14.49 $\pm 0.07$ 16.36 $\pm 0.01$
M/dosage C 0.07 0.15 11.13 11.38 80.19 $\pm 0.08$ 81.74 $\pm 0.17$
M/routeadmin C 0.07 0.11 0.00 0.01 55.23 $\pm 0.15$ 57.59 $\pm 0.09$
M/frequency C 0.08 0.22 12.60 11.31 85.52 $\pm 0.19$ 88.22 $\pm 0.10$
M/prn C 0.00 0.06 0.01 0.08 10.40 $\pm 0.12$ 11.97 $\pm 0.10$